# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v2

@@ Abstract
$\mathcal{R}^{-1}R_{G}=(R_{G}\times\mathcal{R})/\sim\quad\text{where }(\rho_{1},\sigma_{1})\sim(\rho_{2},\sigma_{2})\\ \Leftrightarrow\exists\rho\in\mathcal{R}\text{ such that }\rho\otimes\rho_{1}\otimes\sigma_{2}\simeq\rho\otimes\rho_{2}\otimes\sigma_{1},$ (11)
where $R_{G}$ is the representation ring of $G$, and $\simeq$” is unitary equivalence of representations. Equivalence classes of pairs $(\rho,\sigma)$ are denoted $\rho/\sigma$.
The representation theory of any group which is of relevance to physics is understood in extraordinary detail, so calculating $\mathcal{R}^{-1}R_{G}$, and thus the classification of low-dimensional SALT phases, becomes straightforward. For instance, the classification of $(d,n)=(2,1)$ $U(1)$-SALT phases in Ref. [26] by rational polynomials follows from observing that $R_{U(1)}\cong{\mathbb{Z}}[x,x^{-1}]$ is isomorphic to the ring of integer Laurent polynomials. The fact that including the regular representation, $\rho_{\text{reg}}$, for a finite group in $\mathcal{R}$ leads to a collapse in the classification [93] follows from the fact that $\rho_{\text{reg}}$ is a zero divisor in $R_{G}$. This means Eq. (11) can be satisfied even if $\rho_{1}\otimes\sigma_{2}\not\simeq\rho_{2}\otimes\sigma_{1}$.
@@
Figure 3: Sketch of the pulse schedule for the models defined by the swindle map, with resulting unitary evolution $\tilde{S}$. Sites are coupled along even bonds (red, $t\in[0,\tfrac{1}{3})$) and then along odd bonds (blue, $t\in[\tfrac{1}{3},\tfrac{2}{3})$). Finally, an on-site disorder potential may be applied to promote localization (gray, $t\in[\tfrac{2}{3},1]$).
Models.—While the swindle map [Eq. (8), Fig. 2] was introduced as an abstract tool, it also defines a class of soluble models of $n$-tone-driven (S)ALT phases given any $(n-1)$-parameter family of QCA in one dimension lower. As these models are finite depth circuits, they are also well suited for realization in quantum simulator experiments [63, 64, 65] (Fig. 3).
As an example, Fig. 1(b) illustrates a three-dimensional phase with two driving tones which pumps anomalous Floquet phases to its boundary, and thus has a quantized surface magnetization [52, 27]. This is obtained by applying the swindle map to any model of the anomalous Floquet phase [22, 23].
As three-dimensional systems are typically still challenging for quantum simulators, in Appendix E we construct a soluble model of the quasiperiodic energy pump (QP pump) [31, 32, 53], a one-dimensional two-tone-driven SALT phase with $U(1)$ symmetry. The QP pump supports edge states which pump energy between the two driving fields. The classification of SALT phases shows that the QP pump is stable to strong interactions, provides a nearest-neighbor qubit model for this phase, and indicates that forgetting the $U(1)$ symmetry puts the model in the phase of Refs. [29, 30].
@@
Discussion.—In the context of exactly soluble commuting models and driven phases of localized systems, authors have repeatedly expressed confusion regarding the position of their models within existing classifications [23, 29, 26, 91]. Most of these are classifications of states [66, 115, 28], while localized models (including any commuting projector model) impose nontrivial structure on the entire spectrum. This structure need not be captured by any single state, making such classifications of limited applicability. This point is implicit in the literature [23, 91, 31, 50], but the lack of a systematic and general study seems to have allowed confusion to persist.
@@
The general relationship between (S)ALT phases and QCA we have demonstrated, as well as the $\Omega$-spectrum property which relates static (S)ALT phases to driven phases, extends to fermionic systems essentially without modification (Appendix A-B[56]. The $\Omega$-spectrum structure of the classification can also be exploited to classify SALT phases with lattice symmetries [116, 103, 104]. The more challenging direction of generalization is to include topological order (long-ranged entanglement), where the correspondence with QCA breaks down [20]. This case is of particular interest given recent proposals of using driven topologically ordered localized models to perform quantum computation [117, 118, 119, 120]. The notion of measurement QCA introduced by Ref. [121] may provide a route towards classifying localized phases with topological order, which we will explore in future work.
The authors are particularly grateful to R. Kobayashi for collaboration during the early stages of this project. The authors also thank M. Barkeshli, A. Chandran, L. Fidkowski, S. Pace, R. Thorngren, Y.-T. Tu, B. Ware, and C. Zhang for useful discussions. This work is supported by the Laboratory for Physical Sciences (DML). Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities (DVE).
@@ Appendix A From ALT phases to QCA
We formalize our definitions of “MBL Hamiltonian” and “short-ranged entanglement” thereof in Appendix A.1. The main point of this section and Appendix A.2 is that these physical notions reproduce the formal requirements of [56, Theorem II.4]. Thus, we realize abstract restrictions on commuting projector models as physical properties of localized systems. [56, Theorem II.4] then states that the local integrals of motion (LIOMs) of the MBL Hamiltonian can be prepared by a QCA. We make an equivalent proof in Appendix A.2. The QCA in question is not unique, and we characterize the redundancy in Appendix A.3. We also discuss localized phases of driven systems, which we also characterize (non-uniquely) by parameterized families of QCA.
@@ A.1 Setup
While all classification results in this work are for spin systems on the integer lattice ${\mathbb{Z}}^{d}$, the correspondence between localized phases and QCA holds for either spins or fermions and any locally finite set of sites. That is, any subset of space with finite diameter contains finitely many degrees of freedom. We will call the of points in space supporting some degree of freedom a lattice, though it does not need to have any further regularity.
@@
[where $B_{r}(j)$ is the ball of radius $r$ around $j$] such that $[\tau^{z}_{j},H]=[\tau^{z}_{j},\tau^{z}_{j^{\prime}}]=0$. Further, the set of LIOMs must be maximal, such that a local operator $A$ commutes with all LIOMs only if $A\in\langle\tau^{z}_{j}\rangle_{j\in J}$.
@@
such that $[\tau^{x}_{j},\tau^{z}_{j^{\prime}}]=[\tau^{x}_{j},\tau^{x}_{j^{\prime}}]=0$ for $j\neq j^{\prime}$ and $\tau^{x}_{j}$ cyclically permutes the eigenspaces of $\tau^{z}_{j}$. That is, denoting the ordered set of distinct eigenvalues of $\tau^{z}_{j}$ as $(\lambda^{j}_{0},\ldots\lambda^{j}_{N_{j}-1})$ and the associated eigenspaces $\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})$, we have
@@
This definition should be compared to [56, Definition II.3]. One can also compare this notion of “short-ranged entanglement” for Hamiltonians to the usual notion of short-ranged entanglement for states. A state $|\psi\rangle$ may be said to be short-range entangled if, for any other state $|\psi^{\prime}\rangle$ which differs from $|\psi\rangle$ only in some ball $B_{r}(x)$ [$\langle\psi|A|\psi\rangle=\langle\psi^{\prime}|A|\psi^{\prime}\rangle$ for any local $A$ not supported in $B_{r}(x)$] there is a local unitary $\tau^{x}_{\psi\psi^{\prime}}$ supported in $B_{r}(x)$ such that $\tau^{x}_{\psi\psi^{\prime}}|\psi^{\prime}\rangle=|\psi\rangle$. Any eigenstate of an $H$ satisfying Definition A.2 is short-range entangled in this sense (the required $\tau^{x}_{\psi\psi^{\prime}}$ can be constructed from the $\tau^{x}_{j}$ operators), but the converse—that $\tau^{x}_{j}$ operators exist for any MBL Hamiltonian $H$ whose eigenstates are all short-range entangled—is not obvious.
@@
Here, $B_{r}(S)$ is the ball of radius $r$ around the set $S$. Properly, a QCA should be a projective unitary, and this distinction will be important in later appendices, but not here.
Reference [51] defines a commuting model to be “without intrinsic topological order” when the LIOMs $\tau^{z}_{j}$ can be prepared from single site operators

### v3

@@ IV.2.3 Classification
$\hat{\mathcal{R}}=\{v\in\mathrm{Rep}_{\mathrm{proj}}(G)\,:\,\exists\sigma\in\mathrm{Rep}_{\mathrm{proj}}(G)\\ \text{such that }v\sigma\in\mathcal{R}\}.$ (72)
Indeed, for $v=v_{0}$, this element is $\sigma=v_{1}$, which gives us $\sigma v=\rho_{0}\rho_{1}\in\mathcal{R}$. The unitary similarity transformation relating these is $u_{0,1}$. Note also that $\hat{\mathcal{R}}$ is closed under multiplication by $\mathcal{R}$, so the localization $\mathcal{R}^{-1}\hat{\mathcal{R}}$ is well defined. Happily, $\mathcal{R}^{-1}\hat{\mathcal{R}}$ is actually an abelian group, with $(v/\rho)^{-1}=\rho\sigma/v\sigma$, where the denominator is in $\mathcal{R}$ and the numerator is in $\hat{\mathcal{R}}$ [$v$ demonstrates that there is an element of $\mathrm{Rep}_{\mathrm{proj}}(G)$ such that $v\rho\sigma\in\mathcal{R}$].
All the elements of $\mathcal{R}^{-1}\hat{\mathcal{R}}$ may occur as invariants of some QCA. Suppose $v_{0}/\rho_{0}\in\mathcal{R}^{-1}\hat{\mathcal{R}}$. Then, as $v_{0}$ belongs to $\hat{\mathcal{R}}$, there are representations $v_{1}\in\mathrm{Rep}_{\mathrm{proj}}(G)$ and $\rho_{1}\in\mathcal{R}$ and a unitary $u_{0,1}^{\prime}$ such that
$v_{0}\otimes v_{1}u_{0,1}^{\prime}=u_{0,1}^{\prime}\rho_{1}.$ (73)
We tensor in $\rho_{0}$ to get
where $u_{x,x+1}$ is a translation of $u_{0,1}$ from Eq. (74). The invariant for the QCA $V$ is, by construction, $v_{0}/\rho_{0}$, or its inverse $v_{1}\rho_{0}/\rho_{1}$.
Thus, the full classifying group is
$\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})\cong\log\mathcal{R}^{-1}\hat{\mathcal{R}}$, (76)
where we took a logarithm just to formally make the group operation addition. If $\mathcal{R}=\mathrm{Rep}_{\mathrm{proj}}(G)$ consists of all projective representations of $G$ this becomes the Grothendieck group of $\mathrm{Rep}_{\mathrm{proj}}(G)$, which is by definition
$\pi_{0}(\mathrm{QCA}_{1}^{\mathrm{Rep}_{\mathrm{proj}}(G)})\cong\log\mathrm{Rep}_{\mathrm{proj}}(G)^{-1}\mathrm{Rep}_{\mathrm{proj}}(G).$ (77)
This classifying group can be realized as an extension of a linear part by a projective part. The linear part is a localization of the submonoid of $\mathrm{Rep}_{\mathrm{proj}}(G)$ consisting of projective representations that admit lifts to linear representations. This submonoid is isomorphic to $\mathrm{Rep}(G)/X_{G}$, where $X_{G}$ is the group of units (invertible elements) in $\mathrm{Rep}(G)$—that is, one-dimensional representations—which all become identified as projective representations [67]. The projective part is classified by $H^{2}[G;U(1)]$, as is well known. The induced homomorphism from evaluation on a symmetric product state, $e_{0}^{*}$, just corresponds to sending a projective representation to its classifying cohomology class in $H^{2}[G;U(1)]$, and thus reproduces the group cohomology classification of invertible states in one dimension [71, 66].
The case of $G=U(1)$ serves as a simple example [26, 27]. We take $\mathcal{R}=\mathrm{Rep}_{\mathrm{proj}}(U(1))\cong{\mathbb{N}}[x,x^{-1}]^{\times}/X_{U(1)}$ to consist of all representations of $U(1)$, where the group of units $X_{U(1)}$ consists of all monomials $x^{m}$. (The quotient identifies representations which differ by a global phase.) Then $\hat{\mathcal{R}}=\mathrm{Rep}_{\mathrm{proj}}(U(1))$ and the Grothendieck group is isomorphic to the multiplicative group of rational polynomials with positive coefficients, quotient by the group of units,
$\mathcal{R}^{-1}\hat{\mathcal{R}}\cong\mathrm{Frac}\,{\mathbb{N}}[x]^{\times}/X_{U(1)}.$ (78)
Expressing the quotient in the additive notation, this is
$\log\mathcal{R}^{-1}\hat{\mathcal{R}}\cong(\log\mathrm{Frac}\,{\mathbb{N}}[x]^{\times})/{\mathbb{Z}}\log x.$ (79)
This is the complete classification of $U(1)$-symmetric QCA in one dimension [26].
It has also been observed that including the regular representation for a finite group $G$ in $\mathcal{R}$, $\rho_{\mathrm{reg}}\in\mathcal{R}$, reduces the classification to just $H^{2}[G;U(1)]\oplus\log{\mathbb{Q}}^{\times}$ [66]. Indeed, in the current presentation it is clear that this is because $\mathrm{Rep}(G)$ is not cancellative when $G$ is a finite group. If $\rho_{\mathrm{reg}}$ is the regular representation, then in $\mathrm{Rep}(G)$ we have $\rho_{\mathrm{reg}}\sigma=\mathrm{dim}(\sigma)\rho_{\mathrm{reg}}$, where $\dim(\sigma)$ is the trivial representation with dimension $\dim(\sigma)$. Then
$\rho_{\mathrm{reg}}\sigma\mathrm{dim}(\sigma^{\prime})=\rho_{\mathrm{reg}}\sigma^{\prime}\mathrm{dim}(\sigma),$ (80)
[…]
V Models
The swindle map $S(v,t)$ has been used until now as a theoretical tool. In fact, it also provides a useful construction for explicit models of $d$-dimensional driven (S)ALT phases, given a QCA (or family of QCA) $v$ in one dimension lower. If $v$ is nearest-neighbor, so is $S(v,t)$ (in the $\ell_{\infty}$ distance—it may involve diagonal couplings in the square lattice). Even better, a slightly modified model only involves coupling two qubits at a time in one dimension. Thus, these models are well suited to experimental realization in a variety of quantum simulation platforms [68, 69, 70].
Further, the swindle models are exactly solvable, in the sense that we can predict all observables for all times. Thus, they can explicitly demonstrate what, precisely, is anomalous about ALT phases. Namely, these models demonstrate anomalous edge dynamics—impossible in a short-range entangled MBL model of the same dimension as the edge—as a consequence of pumping nontrivial QCA.
That the swindle construction produces all $n$-tone phases relies on the conjectured $\Omega$-spectrum property, Eq. (9), but the constructed phases being nontrivial and distinct does not.
While the swindle map does not provide models of static ALT phases, we observe that Ref. [110] implies that the edges of static ALT phases also cannot correspond to short-range entangled MBL models. Thus, anomalous edge dynamics are a generic feature of all ALT phases.
In Sec. V.1 we briefly describe a simple class of examples, building on the anomalous Floquet phases of Refs. [22, 23]. These examples are built from swindles applied to a translation QCA, and give a simple picture illustrating the anomalous edge dynamics of ALT phases. Sec. V.2 constructs a more experimentally viable implementation of the swindle map, and Sec. V.3 applies it to a very detailed investigation of a one-dimensional SALT phase with two drives and $U(1)$-symmetry, the quasiperiodic energy pump (QP pump) [31, 32, 53]. We include explicit expressions for the Hamiltonian and quantized observables which reveal the anomalous edge dynamics. The model we obtain through the swindle construction is relatively simple, at least in its operator content, paving the way for experiments on this phase.
V.1 Anomalous edge dynamics
As a basic example, taking $v$ in the swindle $S(v,t)$ to be a one-dimensional translation—or shift—we construct a two-dimensional Floquet phase from a finite depth circuit composed of four layers of $\mathrm{SWAP}$ gates. [A $\mathrm{SWAP}$ gate conjugated by a translation as in Eq. (10) is still a $\mathrm{SWAP}$ gate.] This is precisely the model of the anomalous Floquet phase constructed in Refs. [22, 23], which has circulating bulk orbitals and chiral edge dynamics—each Floquet cycle implements a unidirectional translation $v$ on the left edge (Fig. 2). This behavior is impossible in a localized model in one dimension, both because particles are clearly not localized under this dynamics, and more fundamentally because no Hamiltonian evolution can implement a unidirectional translation.
Iterating the swindle construction on the translation $v$ produces a previously-unidentified phase of three-dimensional two-tone-driven MBL systems, with an explicit finite depth circuit model $S(S(v,t),\nu t)$. This model pumps anomalous Floquet unitaries $S(v,t)$ every period $1/\nu$. To understand this, it is instructive to consider $\nu\ll 1$, which allows us to consider $\nu t$ as a slow adiabatic parameter deforming the period-1 evolution of $S(v,t)$. For $\nu t\approx 0$, the evolution everywhere in the lattice is a small-amplitude perturbation of a trivially localized model. The evolution deforms as $\nu t$ grows, but remains localized. By $\nu t\approx 1$ the evolution in the bulk is again close to trivial, but the evolution of the boundary is now by the anomalous Floquet unitary $S(v,t)$. By $\nu t\approx 2$ the boundary evolution is by $S(v,t)^{2}$—a Floquet evolution with twice the index of $S(v,t)$ in $\pi_{1}(\mathrm{QCA}_{2})$, as witnessed by the circulation of orbitals on the surface [24, 23, 27]. The circulation at the two-dimensional surface continues to grow with time. As the circulation is an invariant of two-dimensional localized Floquet models, this edge evolution is impossible in a strictly two-dimensional localized model.
V.2 Modified swindle
If $v\in\mathrm{QCA}_{d-1}$ has a range $r=1$, then $S(v,t)$ also has range $1$. However, say in one dimension, it typically couples a site $j$ to both j1j-1 and $j+1$ simultaneously. This makes experimental implementation of the swindle models in some quantum simulation platforms more complicated. Further, the swindle model is translationally invariant, and so does not enjoy the robustness of disordered systems to perturbations.
We address both of these issues by defining a modified swindle map $\tilde{S}(v,t)$. We split the loop defining the swindle unitary into three segments, $t\in[0,\tfrac{1}{3})$, $t\in[\tfrac{1}{3},\tfrac{2}{3})$, and $t\in[\tfrac{2}{3},1]$. Within the first segment, sites are coupled along even numbered bonds along the first dimension of the lattice, while odd numbered bonds are coupled in the second segment. The third segment introduces on-site disorder to the model.
@@ V.2 Modified swindle
Figure 4: Sketch of the pulse schedule for the models defined by the modified swindle map $\tilde{S}$. Sites are coupled along even bonds (red, $t\in[0,\tfrac{1}{3})$) and then along odd bonds (blue, $t\in[\tfrac{1}{3},\tfrac{2}{3})$). Finally, an on-site disorder potential may be applied to promote localization (gray, $t\in[\tfrac{2}{3},1]$).
We pick arbitrary smooth functions $\eta_{1,2,3}(t)$ which respectively map $t\in[0,\tfrac{1}{3}]$, $\tfrac{1}{3},\tfrac{2}{3}]$, and $\tfrac{2}{3},1]$ to $\eta\in[0,1]$. Sensible choices for these functions are monotonically increasing. We also choose disordered (d1)(d-1)-dimensional evolution operators $u_{j}^{\mathrm{dis}}(\eta)$, which are required to commute with any desired symmetry of the model. Then, on open boundary conditions with $L\in 2{\mathbb{Z}}$ sites along the first dimension, we have
$\tilde{S}(v,t)=\prod_{j=0}^{L/2-1}u_{2j,2j+1}(v,\eta_{1}(t))$ (81)
[where $u_{j,j+1}$ is defined as in Eq. (10)] for $t\in[0,\tfrac{1}{3})$,
for $t\in[\tfrac{1}{3},\tfrac{2}{3})$, and
$\tilde{S}(v,t)=u_{j}^{\mathrm{dis}}(\eta_{3}(t))v_{0}v^{\dagger}_{L}$ (83)
for $t\in[\tfrac{2}{3},1]$ (Fig. 4).
$H(t)=[i\partial_{t}\tilde{S}(v,t)]\tilde{S}(v,t)^{\dagger}.$ (84)
V.3 QP pump
The model has a $U(1)$ symmetry generated by
$Q=\sum_{j}Q_{j}=\sum_{j}\tfrac{1}{2}(\mathbbm{1}+\sigma^{z}_{j}),$ (85)
$\mathrm{swap}_{j,j+1}(t)=\exp\left[-i\frac{2\pi}{8}t\,\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}\right]$ (86)
$v_{j}(\theta)=\exp\left[-i\frac{W}{2}\theta\sigma^{z}_{j}\right],$ (87)
where $W\in{\mathbb{Z}}$ is the winding number and $\theta\in[0,2\pi)$ parameterizes the loop. All $W\in{\mathbb{Z}}$ give distinct winding numbers with $U(1)$ symmetry. Without this symmetry, all odd (even) winding numbers would be connected by homotopies (Appendix C.3).
For the model to be quasiperiodically driven, we will take $\theta(t)=\omega t\bmod 2\pi$ to be a linear function of $t$ with $\omega\not\in{\mathbb{Q}}$ irrational.
where $\delta_{j}$ is, say, uniformly random in $\Delta,\Delta]$.
The Hamiltonian is then defined piecewise as
where
$h_{j,j+1}^{1}(\eta_{1},\theta)$ $=\frac{W}{4}\left[(1-\cos(\pi\eta_{1}))(\sigma^{z}_{j}-\sigma^{z}_{j+1})-\frac{i}{2}\sin(\pi\eta_{1})e^{-iW\theta}\sigma_{j}^{+}\sigma_{j+1}^{-}+\text{H.c.}\right],\text{ and}$ (92)
$h_{j,j+1}^{2}(\eta_{1},\theta)$ $=\frac{2\pi}{8}\left[\sin(W\theta)\sin(\pi\eta_{1})(\sigma^{z}_{2j}-\sigma^{z}_{2j+1})+\frac{e^{-iW\theta}}{2}(1-\cos(W\theta)-i\sin(W\theta)\cos(\pi\eta_{1}))\sigma_{j}^{+}\sigma_{j+1}^{-}+\text{H.c.}\right].$ (93)
For the second pulse, we have
where
Finally, the disorder pulse has
$H^{\prime}=-\omega\frac{W}{2}\sum_{j=0}^{L-1}j\sigma^{z}_{j}$ (98)
$P(t,\theta)=-\omega\partial_{\theta}H(t,\theta),$ (99)
and $\tau^{z}_{0}=\sigma^{z}_{0}$ otherwise. Equation (101) then gives that
Further using that the power is a sum of traceless local operators, $P=\sum_{j}P_{j,j+1}$, and that $\tau^{z}_{0,L-1}$ have disjoint support, we can expand Eq. (103) as
Calculating the LIOMs from the definition of $u_{0,1}$ gives
The power pumped by the left (smaller $j$) edge state is
VI Discussion
@@
In the context of exactly soluble commuting models and driven phases of localized systems, authors have repeatedly expressed confusion regarding the position of their models within existing classifications [23, 29, 26, 97]. Most of these are classifications of states [71, 119, 28], while localized models (including any commuting projector model) impose nontrivial structure on the entire spectrum. This structure need not be captured by any single state, making such classifications of limited applicability. This point is implicit in the literature [23, 97, 31, 50], but the lack of a systematic and general study seems to have allowed confusion to persist.
@@
Our work demonstrates a general relationship between ALT phases and QCA. This extends both to models with on-site symmetries—forming symmetry-enriched ALT phases, or SALT phases—and to fermionic systems [56] essentially without modification. The $\Omega$-spectrum property we have proposed also allows for the complete classification of all homotopy groups of one-dimensional QCA and SALT phases with any on-site symmetry, reproducing and extending the classifications of Refs. [66, 26, 67].
We expect that the $\Omega$-spectrum structure can also be exploited to classify SALT phases with lattice symmetries [120, 64, 65] (Appendix D). The more challenging direction of generalization is to include topological order (long-ranged entanglement), where the correspondence with QCA breaks down [20]. This case is of particular interest given recent proposals of using driven topologically ordered localized models to perform quantum computation [121, 122, 123, 124, 125].
Beyond the construction of a rigorous proof of the $\Omega$-spectrum conjecture, our work raises several other mathematical questions. For instance, an important extension of our work would be to include a rigorous treatment of tails of LIOMs, as generically arise in any disordered system. This requires working with QCAs which can map local operators to quasilocal operators [110, 126]. We expect that the classification is not altered in any essential way by including tails, but the technical tools used in any proof would be quite different.
The models constructed through the swindle map are finite depth circuits, and are in immediate reach of current quantum simulation experiments [68, 69, 70]. While the three-dimensional geometry of the phase shown in Fig. 2 is difficult to achieve, the model of the QP pump in Sec. V.3 should be straightforward to realize in most popular platforms for quantum simulation. Applying the swindle construction once more to the QP pump model produces a two-dimensional model with three-tone quasiperiodic driving, which has received little prior study [53], in part because exact numerical simulations of many-body quantum dynamics in two dimensions are very challenging. The swindle model provides a solvable point for the associated topological phase, but experimental simulation away from this fixed point may outperform achievable numerics.
Acknowledgements.
The authors are particularly grateful to Ryohei Kobayashi for collaboration during the early stages of this project. The authors also thank Maissam Barkeshli, Anushya Chandran, Lukasz Fidkowski, Jeongwan Haah, Salvatore Pace, Ryan Thorngren, Yi-Ting Tu, Brayden Ware, and Carolyn Zhang for useful discussions. This work is supported by the Laboratory for Physical Sciences, a Stanford Q-FARM Bloch Fellowship (DML), and a Packard Fellowship in Science and Engineering (DML, PI: Vedika Khemani). Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities (DVE).
@@ Appendix A From ALT phases to QCA
We formalize our definitions of “MBL Hamiltonian” and “short-ranged entanglement” thereof in Appendix A.1. The main point of this section and Appendix A.2 is that these physical notions reproduce the formal requirements of [56, Theorem II.4]. Thus, we realize abstract restrictions on commuting projector models as physical properties of localized systems. [56, Theorem II.4] then states that the LIOMs of the MBL Hamiltonian can be prepared by a QCA. We make an equivalent proof in Appendix A.2. The QCA in question is not unique, and we characterize the redundancy in Appendix A.3. We also discuss localized phases of driven systems, which we also characterize (non-uniquely) by parameterized families of QCA.
@@ A.1 Setup
While all classification results in this work are for spin systems on the integer lattice ${\mathbb{Z}}^{d}$, the correspondence between localized phases and QCA holds for either spins or fermions and any locally finite set of sites. That is, any subset of space with finite diameter contains finitely many degrees of freedom. We will call the set of points in space supporting some degree of freedom a lattice, though it does not need to have any further regularity.
@@
[where $B_{r}(j)$ is the ball of radius $r$ around $j$] such that $\tau^{z}_{j},H]=[\tau^{z}_{j},\tau^{z}_{j^{\prime}}]=0$. Further, the set of LIOMs must be complete, such that a local operator $A$ commutes with all LIOMs only if $A\in\langle\tau^{z}_{j}\rangle_{j\in J}$.
@@
such that $\tau^{x}_{j},\tau^{z}_{j^{\prime}}]=[\tau^{x}_{j},\tau^{x}_{j^{\prime}}]=0$ for $j\neq j^{\prime}$ and $\tau^{x}_{j}$ cyclically permutes the eigenspaces of $\tau^{z}_{j}$. That is, denoting the ordered set of distinct eigenvalues of $\tau^{z}_{j}$ as $(\lambda^{j}_{0},\ldots\lambda^{j}_{N_{j}-1})$ and the associated eigenspaces $\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})$, we have
@@
This definition should be compared to [56, Definition II.3]. One can also compare this notion of “short-ranged entanglement” for Hamiltonians to the usual notion of short-ranged entanglement for states. A state $|\psi\rangle$ may be said to be short-range entangled if, for any other state $|\psi^{\prime}\rangle$ which differs from $|\psi\rangle$ only in some ball $B_{r}(x)$ $\langle\psi|A|\psi\rangle=\langle\psi^{\prime}|A|\psi^{\prime}\rangle$ for any local $A$ not supported in $B_{r}(x)$] there is a local unitary $\tau^{x}_{\psi\psi^{\prime}}$ supported in $B_{r}(x)$ such that $\tau^{x}_{\psi\psi^{\prime}}|\psi^{\prime}\rangle=|\psi\rangle$. Any eigenstate of an $H$ satisfying Definition A.2 is short-range entangled in this sense (the required $\tau^{x}_{\psi\psi^{\prime}}$ can be constructed from the $\tau^{x}_{j}$ operators), but the converse—that $\tau^{x}_{j}$ operators exist for any MBL Hamiltonian $H$ whose eigenstates are all short-range entangled—is not obvious.
@@
Here, $B_{r}(S)$ is the ball of radius $r$ around the set $S$. Properly, a QCA should be a projective unitary, but the distinction will not be important here. Reference [51] defines a commuting model to be “without intrinsic topological order” when the LIOMs $\tau^{z}_{j}$ can be prepared from single site operators